OpenStudy (anonymous):

Suppose {a_{1}} > 0 and a_{n+1} = a_{n} + 1/a_{n}. Prove that {a_{n}} diverges to infinity

OpenStudy (anonymous):

Suppose \(a_{1}\ > 0\) and \(a_{n+1}= a_{n} + \frac{1}{a_{n}}\). Prove \(a_{n}\ \rightarrow\ \infty\)

OpenStudy (zarkon):

here is a hint either \(a_1\ge1\) or \(\dfrac{1}{a_1}\ge 1\)

OpenStudy (anonymous):

Couldnt there be a possibility that \(0< a_{n} < 1\)?

OpenStudy (zarkon):

\[a_2=a_1+\frac{1}{a_1}\] \[a_3=a_2+\frac{1}{a_2}\] \[a_4=a_3+\frac{1}{a_3}=a_2+\frac{1}{a_2}+\frac{1}{a_2+\frac{1}{a_2}}>a_2+\frac{1}{a_2}+\frac{1}{a_2+1}\] \[a_5=a_4+\frac{1}{a_4}>a_2+\frac{1}{a_2}+\frac{1}{a_2+1}+\frac{1}{a_2+\frac{1}{a_2}+\frac{1}{a_2+\frac{1}{a_2}}}\] \[>a_2+\frac{1}{a_2}+\frac{1}{a_2+1}+\frac{1}{a_2+1+1}=a_2+\frac{1}{a_2}+\frac{1}{a_2+1}+\frac{1}{a_2+2}\] \[a_n>a_2+\sum_{k=0}^{n-3}\frac{1}{a_2+k}\]

OpenStudy (zarkon):

the last sum, as n goes to infinity, is essentially the harmonic sum. That sum diverges to infinity.

OpenStudy (anonymous):

Wow. Is there anything that just tips you off to do that, or is it just experience?

OpenStudy (zarkon):

well the 1/a1 reminded me on the harmonic. I also used my hint...can you see where?

OpenStudy (anonymous):

Well, I assume it's what allowed you to claim \[a_{2} + \frac{ 1 }{ a_{2} }+ \frac{ 1 }{ a_{2} + \frac{ 1 }{ a_{2} } } > a_{2} + \frac{ 1 }{ a_{2} }+ \frac{ 1 }{ a_{2} + 1 }\] But that means we're saying that its \[\frac{ 1 }{ a_{n} } \ge 1\]Does that mean we need a case where we assume \[a_{n} \ge 1\] or does it work no matter what with that inequality?

OpenStudy (zarkon):

if \(a_1>0\) and either \(a_1\ge 1\) or \(\dfrac{1}{a_1}\ge1\) then \[a_2=a_1+\frac{1}{a_1}>1\] and therefore \(\dfrac{1}{a_2}<1\)

OpenStudy (anonymous):

Oh, I think I see what you mean. So as you continue on with your successive terms, you can say \[\frac{ 1 }{ a_{k} } < 1\] for any k >= 2. At least so it seems. But I see what you did more clearly now. Thank you :)

OpenStudy (zarkon):

yes